# SABR-Informed Multitask Gaussian Process: A Synthetic-to-Real Framework for Implied Volatility Surface Construction

### v1

@@ 3.4 Learning Task Relationships via Hierarchical Regularization
When the target dataset $\mathcal{D}_{\mathcal{T}}$ is much sparser than the source dataset $\mathcal{D}_{\mathcal{S}}$ (our typical scenario), standard maximum marginal likelihood estimation (MLE) can lead to two undesirable outcomes:
@@
$P(\mathbf{e}_{\mathcal{Z}}|\bm{\mu}_{\text{e}},\sigma^{2}_{\text{e}})=\mathcal{N}(\mathbf{e}_{\mathcal{Z}}|\bm{\mu}_{\text{e}},\sigma^{2}_{\text{e}}\mathbf{I}_{d})\quad\text{for }\mathcal{Z}\in\{\mathcal{S},\mathcal{T}\}.$ (17)
Here, $\bm{\mu}_{\text{e}}\in\mathbb{R}^{d^{\prime}}$ and $\sigma^{2}_{\text{e}}$ controlled the squared distance between task embeddings and their mean. We learned both the embeddings $\mathbf{e}_{\mathcal{S}},\mathbf{e}_{\mathcal{T}}$ and the hyperparameters $\bm{\mu}_{\text{e}},\sigma^{2}_{\text{e}}$ from data.
@@
The negative log marginal likelihood $-\log P(\mathbf{y}|\mathbf{X},\bm{\theta})$, derived from the joint Gaussian distribution of observations $\mathbf{y}\sim\mathcal{N}(\mathbf{0},\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma})$ under the MTGP model, is given exactly by:
$-\log P(\mathbf{y}|\mathbf{X},\bm{\theta})=\frac{1}{2}\mathbf{y}^{\intercal}(\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma})^{-1}\mathbf{y}+\frac{1}{2}\log|\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma}|+\frac{M+N}{2}\log(2\pi)$ (20)
@@
Substituting the negative log marginal likelihood (NLML, Equation 20) and the relevant negative log prior (Equation 3.4) into the MAP objective (Equation 18), we arrived at the final objective function $\mathcal{L}(\bm{\theta})$ to minimize:
@@
We optimized Equation 21 using L-BFGS to find:
@@
The optimization balanced fitting the data well (minimizing the NLML term) while keeping task embeddings plausible under their common prior (minimizing the regularization term). The regularization strength adapted based on the learned variance $\sigma^{2}_{\text{e}}$.
@@ 3.5 Prediction
Once we optimized all model parameters, we could make predictions for the target implied volatility surface at any new test point. The prediction followed standard GP principles, adapted for our multi-task framework.
@@
$\begin{bmatrix}\mathbf{y}\\ f_{\mathcal{T}}(\mathbf{x}_{\ast})\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mathbf{0}_{M+N}\\ 0\end{bmatrix},\begin{bmatrix}\mathbf{K}\circ\mathbf{C}+\bm{\Sigma}&\mathbf{k}_{\ast}\\ \mathbf{k}_{\ast}^{\intercal}&C_{\mathcal{T,T}}k(\mathbf{x}_{\ast},\mathbf{x}_{\ast})\end{bmatrix}\right)$ (22)
@@
$\mathbf{k}_{\ast}=\Big{[}\underbrace{C_{\mathcal{T,S}}k(\mathbf{x}_{\ast},\mathbf{x}_{\mathcal{S},1}),\dots,C_{\mathcal{T,S}}k(\mathbf{x}_{\ast},\mathbf{x}_{\mathcal{S},M})}_{\text{Covariance with source (SABR) points}},\underbrace{C_{\mathcal{T,T}}k(\mathbf{x}_{\ast},\mathbf{x}_{\mathcal{T},1}),\dots,C_{\mathcal{T,T}}k(\mathbf{x}_{\ast},\mathbf{x}_{\mathcal{T},N})}_{\text{Covariance with target (market) points}}\Big{]}^{\intercal}$ (23)
@@
4 Experiments Design
To evaluate SABR-MTGP against benchmark methods, we designed numerical experiments. Since the goal was to assess implied volatility surface construction accuracy, we needed a reliable ground truth reflecting realistic market behavior. We chose the Heston stochastic volatility model [12] for this purpose, as it generates complex volatility dynamics similar to real markets.
Our experimental setup included generating realistic ground truth data, creating sparse “market" observations, calibrating the SABR model, training the different models, and evaluating their performance across market conditions. This section details the methodology; Section 5 presents the findings.
@@ 4.1 Data Generation Framework
Our experiments used two main datasets: (1) a sparse “market" dataset simulating real option price observations, and (2) a denser synthetic dataset derived from fitting a SABR model to these observations. The first was the target task, the second was the source task for multi-task learning.
@@ 4.1.1 Heston Ground Truth Data
We generated synthetic “market" data using the Heston model [12], which describes asset price $S_{t}$ and variance $v_{t}$ dynamics via coupled SDEs:
@@
We chose the Heston model because it generates realistic volatility smiles/skews and allows efficient option pricing via its characteristic function for accurate ground truth calculation. Moreover, its parameters have financial interpretations, enabling simulation of various market regimes. For our baseline scenario (the Base case), we used parameters: $v_{0}=0.09$, $\theta=0.09$, $\kappa=1.0$, $\nu_{\text{vol}}=0.8$, and $\rho=-0.8$. We set $S_{0}=100$, $r=0.03$, and $q=0.01$.
@@
To simulate “market" option data, we generated a dataset $\mathcal{D}_{\mathcal{T}}$ containing $N=166$ European call option contracts. We chose the maturities to match typical market distribution. This led to expirations about monthly up to 9 months (0.08 to 0.75 years). Then, we used sparser intervals close to quarterly steps up to 1.75 years. And we included annual expirations at 2.0, 2.5, and 3.0 years. For each maturity, we chose strike prices ($K$) to cover a specific moneyness range ($K/S_{0}$). This range depended on the maturity: $(0.7,1.6)$ for short-term ($\tau\leq 0.5$ years), and $(0.8,1.4)$ for mid-term ($0.5<\tau\leq 1.5$ years) and long-term ($\tau>1.5$ years). Inside these ranges, we changed the strike intervals in a planned way. So, there were more strikes near the at-the-money (ATM) region (from $0.95S_{0}$ to $1.05S_{0}$) and fewer strikes further away. Specifically, the strike intervals around ATM, near ATM, and far from the money were (2.5, 5.0, 15.0) for short-term, (5.0, 10.0, 25.0) for mid-term, and (10.0, 25.0, 50.0) for long-term maturities. Figure 1 showed the data point distribution for our experiments.
@@
Figure 1: Distribution of generated Heston “market" data points (target task $\mathcal{D}_{\mathcal{T}}$) in our experiments, by strike price and time to maturity. Dashed red lines indicate evaluation maturities.
For each contract, we computed the Heston price via Carr-Madan fast Fourier transform method [32]. Then, we obtained the Black-Scholes implied volatility $\sigma_{\text{mkt}}(K,\tau)$ by inverting the Heston price using a root-finding algorithm. This produced the “market" dataset $\mathcal{D}_{\mathcal{T}}=\{(\mathbf{X}_{\mathcal{T}},\mathbf{y}_{\mathcal{T}})\}$, the target task data.
4.1.2 SABR-Derived Synthetic Data
The SABR-MTGP used a denser synthetic dataset derived by fitting the SABR model to the Heston data. This involved several steps. First, for each unique maturity $\tau_{(p)}\in\mathbb{T}_{\text{mkt}}$, SABR parameters ($\alpha_{(p)},\rho_{(p)},\nu_{(p)}$) were calibrated by minimizing MSE between SABR volatilities $\sigma_{\text{SABR}}$ and market volatilities $\sigma_{\text{mkt}}$ from $\mathcal{D}_{\mathcal{T}}$ ($\beta=0.5$ fixed). We used Nelder-Mead optimization with constraints, yielding calibrated parameters $\{(\alpha_{(p)},\rho_{(p)},\nu_{(p)})\}$ for the set $\mathbb{T}_{\text{mkt}}$.
Second, we obtained SABR parameters for any time-to-maturity $\tilde{\tau}$ via piecewise linear interpolation (Equation 9) between parameters at the nearest surrounding calibrated maturities in $\mathbb{T}_{\text{mkt}}$. Constant extrapolation was used outside this range. Figure 2 shows the calibrated SABR parameters across maturities for the Base Heston scenario. The parameters showed term structures consistent with market behavior, supporting the interpolation.
@@ 4.1.2 SABR-Derived Synthetic Data
Third, we defined a dense $35\times 35$ grid $(\tilde{K}_{j},\tilde{\tau}_{j})$, covering strike range ($K\in[70,160]$) and maturity range ( $\tau\in[0.08,3.0]$). At each grid point, we used the interpolated/extrapolated SABR parameters with the Hagan formula (Equation 7) to compute synthetic implied volatility $y_{\mathcal{S},j}=\sigma_{\text{SABR}}(\tilde{K}_{j},F_{\tilde{\tau}_{j}},\tilde{\tau}_{j};\alpha_{\tilde{\tau}_{j}},\beta,\rho_{{\tilde{\tau}}_{j}},\nu_{{\tilde{\tau}}_{j}}).$
@@ 4.2 Comparative Models and Training
We compared three approaches for constructing the IVS. First, SABR Interpolation, a parametric baseline using calibrated and interpolated SABR parameters with the Hagan formula (Equation 7). Second, a Single-Task Gaussian Process (GP) trained only on the sparse Heston “market" data $\mathcal{D}_{\mathcal{T}}$. This GP used a Matérn 5/2 kernel on $(K,\tau)$ inputs and learned hyperparameters via L-BFGS optimization of marginal likelihood. Third, the proposed SABR-Informed Multi-Task GP (SABR-MTGP), our transfer learning framework. It was trained jointly on synthetic source data $\mathcal{D}_{\mathcal{S}}$ and market target data $\mathcal{D}_{\mathcal{T}}$, using the ICM structure with a shared Matérn 5/2 input kernel and a task kernel based on learned one-dimensional embeddings.
@@ 4.3 Evaluation Framework and Robustness Analysis
We evaluated model performance against the Heston ground truth on a predefined evaluation grid for fair comparison. This grid, separate from training data, included points $(K_{\ast},\tau_{\ast})$ spanning a relevant moneyness range (e.g., $K_{\ast}/S_{0}\in[0.8,1.4]$) at specific test maturities $\tau_{\ast}$ (e.g., 0.3, 0.9, 2.2 years) representing near-, mid-, and long-term. For each evaluation point, we computed the ground truth Heston IV using a high-precision FFT approach with the Heston characteristic function, followed by numerical inversion to obtain corresponding implied volatility.
Each model (SABR Interpolation, GP, SABR-MTGP) predicted $\hat{\sigma}(K_{\ast},\tau_{\ast})$ on this grid. We measured performance using Root Mean Squared Error (RMSE) and Mean Absolute Error (MeanAE) against the ground truth $\sigma_{\text{Heston}}(K_{\ast},\tau_{\ast})$, calculated over valid predictions.
To assess robustness, we repeated experiments using ten Heston parameter configurations (Table 1), including the Base scenario. These settings simulated various market dynamics by altering parameters like mean reversion $\kappa$, volatility levels $\theta,\nu_{\text{vol}}$, correlation $\rho$, and initial term structure $v_{0}$ vs $\theta$. This allowed analysis across different market conditions. Robustness evaluation focused on metrics at distinct test maturities. We also analyzed the MTGP’s learned task correlation and variance decomposition for each scenario to understand its adaptive information-sharing.
@@ 5 Results and Discussion
This section presents our findings on SABR-MTGP performance compared with SABR interpolation and the standard GP. We first analyze performance under the Base Heston scenario, then examine robustness across various market conditions.
5.1 Performance in the Base Scenario
Figures 3(a), 3(b), and 3(c) show fitted implied volatility curves from the three models compared against Heston ground truth, for near-term ($\tau=0.30$), mid-term ($\tau=0.90$), and long-term ($\tau=2.20$) maturities in the Base scenario. Each figure includes fitted curves and residuals.
(a) Near-term ($\tau=0.30$).
(b) Mid-term ($\tau=0.90$).
(c) Long-term ($\tau=2.20$).

### v2

@@ 3.4 Learning Task Relationships via Hierarchical Regularization
When the target dataset $\mathcal{D}_{\mathcal{T}}$ is much sparser than the source dataset $\mathcal{D}_{\mathcal{S}}$, standard maximum marginal likelihood estimation (MLE) can lead to two undesirable outcomes:
@@
[…]
Here, $\bm{\mu}_{\text{e}}\in\mathbb{R}^{d^{\prime}}$ and $\sigma^{2}_{\text{e}}$ controls the squared distance between task embeddings and their mean. We learned both the embeddings $\mathbf{e}_{\mathcal{S}},\mathbf{e}_{\mathcal{T}}$ and the hyperparameters $\bm{\mu}_{\text{e}},\sigma^{2}_{\text{e}}$ from data.
@@
The negative log marginal likelihood $\log P(\mathbf{y}|\mathbf{X},\bm{\theta})$, derived from the joint Gaussian distribution of observations $\mathbf{y}\sim\mathcal{N}(\mathbf{0},\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma})$ under the MTGP model, is given exactly by:
$\log P(\mathbf{y}|\mathbf{X},\bm{\theta})=\frac{1}{2}\mathbf{y}^{\intercal}(\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma})^{-1}\mathbf{y}+\frac{1}{2}\log|\mathbf{K}\circ\mathbf{C}+\mathbf{\Sigma}|+\frac{M+N}{2}\log(2\pi)$ (18)
@@
Substituting the negative log marginal likelihood (NLML, equation (18)) and the relevant negative log prior (equation (3.4)) into the MAP objective (equation (16)), we arrived at the final objective function $\mathcal{L}(\bm{\theta})$ to minimize:
@@
We optimized equation (19) using L-BFGS to find:
@@
The optimization balanced fitting the data well while keeping task embeddings plausible under their common prior. The regularization strength adapted based on the learned variance $\sigma^{2}_{\text{e}}$.
@@ 3.5 Prediction
Once we optimized all model parameters, we could make predictions for the target implied volatility surface at any new test point. The prediction followed standard GP principles, adapted for our multitask framework.
@@
[…]
@@
[…]
@@
4 Experimental Design
To evaluate SABR-MTGP against benchmark methods, we designed numerical experiments. Since the goal was to assess implied volatility surface construction accuracy, we required a reliable ground truth reflecting realistic market behavior. We chose the Heston stochastic volatility model [25] for this purpose, as it generates complex volatility dynamics similar to real markets.
Our experimental setup comprised the generation of ground-truth data, construction of sparse “market” observations, slice-wise SABR calibration, model training, and evaluation across market regimes. This section details the methodology, and Section 5 presents the findings.
@@ 4.1 Data Generation Framework
We used two datasets: a sparse “market” dataset and a denser synthetic dataset obtained by SABR calibration on the market dataset. The market dataset serves as the target task, and the synthetic dataset serves as the source task for multitask learning.
@@ 4.1.1 Heston Ground Truth Data
We generated synthetic “market” data using the Heston model [25], which describes asset price $S_{t}$ and variance $v_{t}$ dynamics via SDEs:
@@
We chose Heston for its ability to reproduce realistic smiles and skews. Its characteristic function based pricing enables high-precision computations for ground-truth generation. Moreover, its parameters have financial interpretations, enabling simulation of various market regimes. For our baseline scenario (the Base case), we used parameters: $v_{0}=0.09$, $\theta=0.09$, $\kappa=1.0$, $\nu_{\text{vol}}=0.8$, and $\rho=-0.8$. We set $S_{0}=100$, $r=0.03$, and $q=0.01$.
@@
To simulate “market” option data, we generated a dataset $\mathcal{D}_{\mathcal{T}}$ comprising $N=166$ European call option contracts. We designed maturities to mirror market availability, which is dense at short horizons and sparser at longer horizons. This led to expirations about monthly up to 9 months (0.08 to 0.75 years). Then, we used sparser intervals close to quarterly steps up to 1.75 years. Additionally, we included annual expirations at 2.0, 2.5, and 3.0 years. For each maturity, we chose strike prices ($K$) to cover a specific moneyness range ($K/S_{0}$). This range depended on the maturity: $[0.7,1.6]$ for short-term ($\tau\leq 0.5$ years), and $[0.8,1.4]$ for mid-term ($0.5<\tau\leq 1.5$ years) and long-term ($\tau>1.5$ years). Inside these ranges, we changed the strike intervals in a planned way. Consequently, there were more strikes near the at-the-money (ATM) region (from $0.95S_{0}$ to $1.05S_{0}$) and fewer strikes further away. Specifically, the strike intervals around ATM, near ATM, and far from the money were (2.5, 5.0, 15.0) for short-term, (5.0, 10.0, 25.0) for mid-term, and (10.0, 25.0, 50.0) for long-term maturities. Figure 1 shows the data point distribution for our experiments.
@@
Figure 1: Distribution of generated Heston “market” data points (target task $\mathcal{D}_{\mathcal{T}}$) in our experiments, by strike price and time to maturity. Dashed red lines indicate evaluation maturities.
For each option, we computed the Heston price via the Carr-Madan fast Fourier transform method (FFT) [6] using standard damping $\alpha=2.5$ and a high-resolution grid with $N=2^{21}$ and $c=12800$. We verified numerical stability by doubling $N$ and $c$, which changed call prices by less than $10^{-6}$. Here, $N$ denotes the number of frequency samples (FFT grid size) and $c$ is the integration-range truncation parameter in Carr-Madan FFT. The spacing $\eta=c/N$ and the log-strike span $\Lambda=\tfrac{2\pi}{N\eta}$ in the Carr-Madan formulation control the tradeoff between aliasing and accuracy. IVs are computed via a safeguarded Newton method (tolerance $10^{-5}$, maximum 200 iterations, $\sigma\in[10^{-4},5]$). This produced the “market” dataset $\mathcal{D}_{\mathcal{T}}=\{(\mathbf{X}_{\mathcal{T}},\mathbf{y}_{\mathcal{T}})\}$, the target task data.
4.1.2 SABR-generated Synthetic Data
The SABR-MTGP used a denser synthetic dataset derived by fitting the SABR model to the Heston data. We calibrated the SABR following the scheme in Section 3.2. Figure 2 shows the calibrated SABR parameters across maturities for the Base Heston scenario.
@@ 4.1.2 SABR-generated Synthetic Data
Then, we defined a dense $35\times 35$ grid $(\tilde{K}_{j},\tilde{\tau}_{j})$, covering strike range ($K\in[70,160]$) and maturity range ($\tau\in[0.08,3.0]$). At each grid point, we used the Hagan formula equation (5) to compute synthetic implied volatility $y_{\mathcal{S},j}=\sigma_{\text{SABR}}(\tilde{K}_{j},F_{\tilde{\tau}_{j}},\tilde{\tau}_{j}).$
@@ 4.2 Comparative Models and Training
We evaluated six methods: (i) SABR [^1] Hagan asymptotics with fixed $\beta=0.5$; slice-wise calibration of $(\alpha,\rho,\nu)$., (ii) Single-Task GP [^2] Matérn 5/2 kernel on $(K,\tau)$ with L-BFGS training, (iii) the proposed SABR-Informed Multitask GP (SABR-MTGP) [^3] ICM structure with a shared Matérn 5/2 input kernel and a task kernel based on learned one-dimensional embeddings; optimized by L-BFGS., (iv) arbitrage-free SSVI [17], (v) Cubic Spline, and (vi) Neural Network with SABR pre-training [^4] Two-hidden-layer MLP (64 units each, ReLU) pre-trained on synthetic SABR data and fine-tuned on target data with Adam optimizer..
@@ 4.3 Evaluation Framework and Robustness Analysis
We evaluated model performance against the Heston ground truth on a predefined evaluation grid for a fair comparison. This grid, separate from training data, included points $(K_{\ast},\tau_{\ast})$ spanning a relevant moneyness range (e.g., $K_{\ast}/S_{0}\in[0.8,1.4]$) at specific test maturities $\tau_{\ast}$ (e.g., 0.3, 0.9, 2.2 years) representing near-, mid-, and long-term. For each point, we computed Heston prices via the Carr-Madan FFT and inverted to IVs by safeguarded Newton.
Each model (SABR, GP, SABR-MTGP, SSVI, Cubic Spline, and NN) predicted $\hat{\sigma}(K_{\ast},\tau_{\ast})$ on this grid. We measured performance using Root Mean Squared Error (RMSE) and Mean Absolute Error (MAE) against the ground truth $\sigma_{\text{Heston}}(K_{\ast},\tau_{\ast})$, calculated over valid predictions.
To assess robustness, we repeated experiments using ten Heston parameter configurations (Table 1), including the Base scenario. These settings simulated various market dynamics by altering parameters like mean reversion $\kappa$, volatility levels $\theta,\nu_{\text{vol}}$, correlation $\rho$, and initial term structure $v_{0}$ vs $\theta$.
@@ 5 Results and Discussion
This section reports cross regime robustness and then uses Base scenario slices as diagnostics. We compare SABR, the single-task GP, SABR-MTGP, SSVI, cubic spline, and a neural network baseline.
5.1 Performance and Robustness Across Market Conditions
Tables 2 and 3 summarize model accuracy (RMSE and MAE) across ten Heston settings at near-term and long-term horizons. Values are scaled by $10^{-3}$, and the best metric value per setting is shown in bold. At the short maturity, SABR-MTGP is best in eight settings and GP leads in two. At the long maturity, the leader alternates by regime, with SABR best in six settings, SABR-MTGP best in three, and GP best in one. SSVI and cubic spline are competitive in a few settings yet they are not consistently dominant. The neural network baseline is generally weaker.
Table 2: Error summary at Near-term maturity ($\tau=0.3$). Values are scaled ($\times 10^{-3}$).
Setting SABR GP SABR-MTGP SSVI Cubic Spline NN
RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
Base 5.78 5.22 1.93 1.63 1.69 1.44 4.27 3.42 2.11 1.80 4.67 3.49
Moderate Mean-Rev 4.50 4.04 1.55 1.29 1.42 1.25 4.09 3.17 1.79 1.54 6.31 4.36
Low Mean-Rev 6.60 5.97 2.24 1.79 2.09 1.83 4.34 3.39 2.28 1.92 4.91 3.75
High Vol-Regime 4.99 4.56 0.83 0.69 0.71 0.61 3.17 2.50 1.44 1.10 2.67 1.98
Low Vol-Regime 6.74 6.36 4.15 2.48 3.94 2.35 4.04 3.15 5.04 3.17 10.89 7.37
Moderate Correlation 4.10 3.72 1.38 1.08 1.23 1.05 1.90 1.71 1.93 1.45 3.21 2.66
Strong Correlation 7.26 6.13 4.28 3.25 3.10 2.41 7.04 5.44 3.84 3.03 11.51 7.41
Term Structure Up 4.88 4.37 1.34 1.14 1.59 1.41 4.42 3.34 1.84 1.57 5.02 3.60
Term Structure Down 5.24 4.76 1.12 0.96 0.82 0.72 6.25 4.75 1.43 1.09 3.57 2.43
Mixed Regime 2.16 2.00 0.38 0.29 0.89 0.65 1.12 0.83 0.75 0.56 2.03 1.52
Table 3: Error summary at Long-term maturity ($\tau=2.2$). Values are scaled ($\times 10^{-3}$).
Setting SABR GP SABR-MTGP SSVI Cubic Spline NN
RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE RMSE MAE
Base 0.42 0.27 0.82 0.69 0.61 0.59 0.45 0.33 1.08 0.66 4.60 4.10
Moderate Mean-Rev 0.36 0.26 1.18 0.99 0.86 0.72 0.83 0.73 0.39 0.27 3.44 2.87
Low Mean-Rev 0.55 0.45 1.60 1.17 1.81 1.56 1.60 1.32 3.15 1.96 5.24 4.48
High Vol-Regime 0.33 0.24 0.24 0.19 0.17 0.14 0.90 0.76 0.43 0.31 4.05 3.53
Low Vol-Regime 0.51 0.32 2.52 1.73 0.38 0.33 0.75 0.59 1.17 0.72 3.66 2.90
Moderate Correlation 0.12 0.10 1.24 1.19 0.35 0.27 0.65 0.59 1.18 0.79 4.99 4.30
Strong Correlation 0.74 0.41 1.36 0.67 0.15 0.12 0.83 0.74 0.58 0.34 4.01 3.54
Term Structure Up 0.42 0.29 0.32 0.20 0.95 0.83 0.49 0.40 0.49 0.32 3.92 3.25
Term Structure Down 0.57 0.54 0.97 0.84 0.68 0.47 2.34 1.87 1.69 1.22 3.66 3.24
Mixed Regime 0.10 0.07 0.49 0.46 0.13 0.11 0.40 0.34 0.34 0.25 4.53 3.96
Figures 3 and 4 provide visual evidence from the Base scenario. In the near-term slice, SABR shows a systematic error pattern relative to the Heston truth that is consistent with the limitations of a rigid parametric form. The single-task GP follows the curve yet exhibits small local oscillations in the wings where observations are sparse. SABR-MTGP tracks the truth closely and its residuals are centered and smooth.